-7^2 , (-7)^2 why are they different?
Answer by stanbon(75887)

-7^2
The negative is excluded from the squaring:
-7^2 = -7*7 = -49
----
(-7)^2
The negative is included in the squaring:
(-7)^2 = (-7)(-7) = +49
===========================
Cheers,
Stan H.
===========